OpenStudy (anonymous):

Simplify. (343)1/3 A.) 21 B.) 49 C.) 7 D.) 19 My guess?: 7.

OpenStudy (pinklion23):

343/0.3333333333

OpenStudy (anonymous):

i think 7 is right is \(7^3=343\)?

OpenStudy (pinklion23):

oops sorry i meant 0.333333*343

OpenStudy (anonymous):

yeah \(7^3=343\) so \(\sqrt[3]{343}=7\)

OpenStudy (anonymous):

It was 7. Thank you all :)